OpenStudy (anonymous):

Multiply (√5+1)(8√5+1)

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

answer was 41+9√5 no idea how it got there

OpenStudy (skullpatrol):

Each of the numerical expressions within the parentheses is called a binomial, do you recall how to use the FOIL method of multiplying binomials?

OpenStudy (anonymous):

was it x^2+2xy+y^2 for this one?

OpenStudy (anonymous):

ah nvm, as for Foil i need steps to show me for this one for others i know what to do

OpenStudy (skullpatrol):

1. Multiply the first terms of the binomials

OpenStudy (anonymous):

5*8√5*√5+8√5+1 looks wrong but eh

OpenStudy (skullpatrol):

\[\sqrt{5}*8 \sqrt{5} = ?\]

OpenStudy (anonymous):

for the first step?

OpenStudy (skullpatrol):

Yes

OpenStudy (anonymous):

i know where 41 comes from but no idea where 9√5 comes in

OpenStudy (anonymous):

8*5 =40+1 = 41 right?

OpenStudy (anonymous):

can you like show me all the steps how its done so i can duplicate that method and learn how it was done?

OpenStudy (skullpatrol):

\[\sqrt{5}*8\sqrt{5} = 8\sqrt{25}=8*5=40\]

OpenStudy (anonymous):

√5 *√5 = 5, you know that's how square root work, right?

OpenStudy (anonymous):

alright i get that so far

OpenStudy (skullpatrol):

2. Multiply the first term of each binomial by the last term of the other and add these products.

OpenStudy (anonymous):

so would the second line be 8√5+1?

OpenStudy (anonymous):

40+8√5+1?

OpenStudy (skullpatrol):